# Rusty Linux: Advances in Rust for Linux Kernel Development

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: General and reference Surveys and overviewsccs: Software and its engineering Language featuresccs: Computer systems organization Architectures
@@ 2. Research Methodology
For our research methodology, we followed the Kitchenham and Charters (Kitchenham, 2004) methodology to conduct SLR in software engineering. We divided our review into three discrete phases, planning the review, conducting the review, and reporting the review results. The following sections detail our review process and are diagrammed in figure 1.
@@ 2.1. Planning
To begin our study, we must first confirm the need for a SLR. The recent report released by ONCD (United States Gov, 2024) has conveniently done this job for us by compiling a report detailing the need for research in the domain of memory safety. While the ONCD report detailed a two-pronged approach, for this reflection paper, we will be focusing on a memory-safe programming language, specifically Rust. While there are many modern memory-safe programming languages available for software developers to use, Rust is one of the few languages that is feasible to use when developing operating system kernels due to its lack of runtime and garbage collector (Ojeda, 2021). The Linux kernel is currently in the early stages of adding real support for Rust (The kernel development community, [n. d.]) thus, we will focus on Rust as a primary candidate to replace the aging C programming language.
@@ 2.3. Data Collection
ACM (Abstract:(”operating system” kernel linux OS) AND Abstract:(rust) )OR (Title:(”operating system” OS kernel linux) AND Title:(rust)) 30
@@
Monolithic (The kernel development community, [n. d.]), (Li et al., 2019), (Miller et al., 2021), (Oikawa, 2023) Linux kernel v6.1+
@@ 2.4. Data Analysis
Once we had our initial collection of papers from the databases, we merged all duplicate papers into one record and then started our first pass, which consisted of reading the title and abstract and either marking the paper as include or exclude. Once we completed the initial screening, we were left with 26 papers that needed an in-depth reading and analysis. We found an additional 6 papers during forward and backward snowballing that were also reviewed, bringing the total number of papers to 32. The first author manually reviewed all 32 papers, built the first dataset, and conducted the analysis. The second author then verified all the data and results. Since the results were straightforward and came from a comparatively small number of papers, there were no discrepancies or missing results to solve.
@@ 3.1. RQ1: Existing Approaches
The integration of Rust into operating system kernel development has garnered significant attention due to Rust’s promise of memory safety and concurrency guarantees. Several approaches and methodologies have emerged to leverage Rust for kernel development, each addressing different aspects of kernel functionality and integration. We found several approaches that researchers are taking to integrate Rust into kernel development. The first approach is a greenfield approach where the operating system architecture is redesigned and built from the ground up in order to fully leverage the Rust programming language (Culic et al., 2022; Boos et al., 2020). The second approach is the incremental integration of Rust into existing kernels, such as the Linux kernel (The kernel development community, [n. d.]; Miller et al., 2021; Oikawa, 2023). This methodology involves gradually rewriting or augmenting specific components of the kernel with Rust while maintaining compatibility with the existing C codebase. Finally, researches are porting existing C kernels over to Rust and then evaluating the differences. Table 2 summarizes our findings regarding what existing approaches researchers are using to integrate rust into the kernel.
@@
For the second approach, researchers take a more traditional micro-kernel implementation (Chen et al., 2023; Liang et al., 2021). By pushing most kernel functionality to user space, the authors limited the surface area that needs to be proved. Chen X et al. (Chen et al., 2023) took their implementation one step further by formally verifying their new micro-kernel named Atmosphere by using both the liner type system of Rust in combination with an SMT solver. They were able to get a 7.5:1 proof-to-code ratio, which is higher than other formerly verified micro-kernels SeL4 and CeriKOS both have proof code ratios of 19:1 and 20:1, respectively. While this approach was novel drivers running in userspace are still not formally verified and don’t have to be written in Rust.
@@ 3.3. RQ3: Challenges and Limitations
Table 4. Challenges Unique to the Rust Programming Language

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 18th ACM / IEEE International Symposium on Empirical Software Engineering and Measurement; October 24–25, 2024; Barcelona, Spainbooktitle: Proceedings of the 18th ACM / IEEE International Symposium on Empirical Software Engineering and Measurement (ESEM ’24), October 24–25, 2024, Barcelona, Spaindoi: 10.1145/3674805.3690756isbn: 979-8-4007-1047-6/24/10ccs: General and reference Surveys and overviewsccs: Security and privacy Operating systems securityccs: Software and its engineering Operating systemsccs: Software and its engineering Language features
@@ 2. Research Methodology
For our research methodology, we followed the Kitchenham and Charters (Kitchenham, 2004) methodology to conduct a SLR in software engineering. We divided our review into three discrete phases, planning the review, conducting the review, and reporting the review results. The following sections detail our review process and are diagrammed in figure 1.
@@ 2.1. Planning
To begin our study, we must first confirm the need for a SLR. The recent report released by ONCD (United States Gov, 2024) has conveniently done this job for us by compiling a report detailing the need for research in the domain of memory safety. While the ONCD report detailed a two-pronged approach, for this reflection paper, we will be focusing on a memory-safe programming language, specifically Rust. While there are many modern memory-safe programming languages available for software developers to use, Rust is one of the few languages that is feasible to use when developing operating system kernels due to its lack of runtime and garbage collector (Ojeda, 2021). The Linux kernel is currently in the early stages of adding real support for Rust (The kernel development community, 2024) thus, we will focus on Rust as a primary candidate to replace the aging C programming language.
@@ 2.3. Data Collection
ACM (Abstract:(”operating system” OR Kernel OR linux OR OS) AND Abstract:(Rust)) OR (Title:(”operating system” OR OS OR kernel OR linux) AND Title:(Rust)) 30
@@
Monolithic (The kernel development community, 2024), (Li et al., 2019), (Miller et al., 2021), (Oikawa, 2023) Linux kernel v6.1+
@@ 2.4. Data Analysis
Once we had our initial collection of papers from the databases, we merged all duplicate papers into one record and then started our first pass, which consisted of reading the title and abstract and either marking the paper as include or exclude. Once we completed the initial screening, we were left with 22 papers that needed an in-depth reading and analysis. We found an additional 6 papers during forward and backward snowballing that were also reviewed, bringing the total number of papers to 28. The first author manually reviewed all 28 papers, built the first dataset, and conducted the analysis. The second author then verified all the data and results. Since the results were straightforward and came from a comparatively small number of papers, there were no discrepancies or missing results to solve.
@@ 3.1. RQ1: Existing Approaches
The integration of Rust into operating system kernel development has garnered significant attention due to Rust’s promise of memory safety and concurrency guarantees. Several approaches and methodologies have emerged to leverage Rust for kernel development, each addressing different aspects of kernel functionality and integration. We found several approaches that researchers are taking to integrate Rust into kernel development. The first approach is a greenfield approach where the operating system architecture is redesigned and built from the ground up in order to fully leverage the Rust programming language (Culic et al., 2022; Boos et al., 2020). The second approach is the incremental integration of Rust into existing kernels, such as the Linux kernel (The kernel development community, 2024; Miller et al., 2021; Oikawa, 2023). This methodology involves gradually rewriting or augmenting specific components of the kernel with Rust while maintaining compatibility with the existing C codebase. Finally, researches are porting existing C kernels over to Rust and then evaluating the differences. Table 2 summarizes our findings regarding what existing approaches researchers are using to integrate rust into the kernel.
@@
For the second approach, researchers take a more traditional micro-kernel implementation (Chen et al., 2023; Liang et al., 2021). By pushing most kernel functionality to user space, the authors limited the surface area that needs to be proved. Chen X et al. (Chen et al., 2023) took their implementation one step further by formally verifying their new micro-kernel named Atmosphere by using both the liner type system of Rust in combination with an SMT solver. They were able to get a 7.5:1 proof-to-code ratio, which is higher than other formerly verified micro-kernels SeL4 and CeriKOS both of which have proof code ratios of 19:1 and 20:1, respectively. While this approach was novel, drivers running in userspace are still not formally verified and don’t have to be written in Rust.
@@ 3.3. RQ3: Challenges and Limitations
Table 4. Challenges Unique to the Rust Programming Language